mathslover (mathslover):

Logarithm Question Solve \(x + \log_{10} (1+2^x) = x\log_{10} 5 + \log_{10} 6\)

mathslover (mathslover):

The first step I followed was : \(\log_{10} ({10}^x) + \log_{10} ( 1 + 2^x) = x\log_{10} 5 + \log_{10} 6\)

mathslover (mathslover):

Thus, LHS becomes : \(\log_{10} ((10^x)(1+2^x)) = x\log_{10} 5 + \log_{10} 6\)

mathslover (mathslover):

Thus, I get : \((10^x)(1+2^x) = (5^x)(6) \)

mathslover (mathslover):

\((2^x)(1+2^x) = 6\) \(\implies 2^x + 2^{2x} = 6\) -------------- (1) Let \(2^x = y\) Equation (1) becomes : \(y + y^2 - 6 =0 \) or \(y^2 + y - 6 =0 \)

mathslover (mathslover):

Okay, so, using Quadratic Formula now : \(y = \cfrac{-1 \pm \sqrt{1 + 24}}{2} = \cfrac{-1 \pm 5}{2} \) \(\text{Thus y } \space = 2 \) or \(-3\) Since, \(2^x \ne -3\) So, \(2^x = 2\) Thus, x = 1.

mathslover (mathslover):

o.O Made it more like a tutorial :P Yeah, well, I didn't know that I was on a right track but when I started writing that all, I just kept going... Though, I will appreciate your's suggestions and any short methods (if available).

OpenStudy (kainui):

Yeah, kinda cool how often it happens like this. If you see a problem that you know has an answer and you just start walking through it and believe in the math, then you often times just finish it without barely knowing haha

mathslover (mathslover):

Yeah, that's true! :)